# On the connection between Bochner’s theorem on positive definite maps and Choi theorem on complete positivity

### v1

@@ III.5 Relation between complete positivity of maps and positivity of their Fourier transforms
The CP vs. Positivity Correspondence between a linear map $\Phi:M_{m}(\mathbb{C})\to M_{n}(\mathbb{C})$ and its Fourier transform $\widehat{\Phi}(\rho)$ with respect to a representation $\rho:M_{m}\to M_{d_{\rho}}(\mathbb{C})$ holds if and only if $\rho$ is of the form $\rho(X)=UXU^{\dagger}$ for all $X\in M_{m}(\mathbb{C})$, where $U:\mathbb{C}^{m}\to\mathbb{C}^{d_{\rho}}$ is unitary with $d_{\rho}=m$.
@@
$=(\mathrm{id}\otimes\Phi)\sum_{i,j}\rho(e_{ij})\otimes e_{ij}.$
Now it can be shown that $\sum_{i,j}\rho(e_{ij})\otimes e_{ij}=\tau(C_{\rho^{*}})$, where $\rho^{*}$ is the adjoint of the representation $\rho:M_{m}\to M_{d_{\rho}}(\mathbb{C})$, $C_{\rho^{*}}$ is the Choi matrix of $\rho^{*}$ and $\tau$ represents transposition operation. With this relation, the above equation becomes
$\widehat{\Phi}(\rho)$ $=(\mathrm{id}\otimes\Phi)\tau(C_{\rho^{*}})=(\tau\otimes\Phi\circ\tau)C_{\rho^{*}}$
$\iff\tau(\widehat{\Phi}(\rho))=(\mathrm{id}\otimes\tau\circ\Phi\circ\tau)C_{\rho^{*}}.$
@@
As $\tau(\widehat{\Phi}(\rho))$ is positive semidefinite if and only if $\widehat{\Phi}(\rho)$ is positive semidefinite, and $\Phi$ is CP if and only if $\tau\circ\Phi\circ\tau$ is CP, then by using Theorem (1) of [21] (Also see [22]), we have that the CP vs positivity correspondence holds between $\Phi$ and $\widehat{\Phi}(\rho)$ if and only if $\rho^{*}$ is a complete order isomorphism if and only if $\rho$ is a complete order isomorphism. Since $\rho$ is a representation, it must be of the form $\rho(X)=UXU^{\dagger}$ for all $X\in M_{m}(\mathbb{C})$, where $U:\mathbb{C}^{m}\to\mathbb{C}^{d_{\rho}}$ is unitary with $d_{\rho}=m$. This completes the proof. $\Box$
IV Conclusion
The Choi-Jamiołkowski isomorphism is an essential tool in the study of various mathematical objects in quantum information theory. In this work, we have defined convolution product between maps from the contracted semigroup algebra $\mathbb{C}_{0}[S]$ of a semigroup $S$ to an arbitrary associative algebra $\mathcal{A}$. The convolution product makes the space $L(\mathbb{C}_{0}[S],\mathcal{A})$ of linear maps from $\mathbb{C}_{0}[S]$ to $\mathcal{A}$ an associative algebra. Then we have proved that the convolution algebra $L(\mathbb{C}_{0}[S],\mathcal{A})$ and the tensor product algebra $\mathbb{C}_{0}[S]\otimes\mathcal{A}$ are isomorphic. In the specific case of matrix unit semigroup, our result shows that the product in the space of maps on the matrix algebras which is preserved by the Choi-Jamiołkowski isomorphism is convolution. Then we defined Fourier transform of a map from $\mathbb{C}_{0}[S]$ to $M_{n}(\mathbb{C})$ and derived the Fourier inversion formula when $S$ is a finite inverse semigroup. In the case of the inverse semigroup of matrix units, the Fourier transformation of a map with respect to the identity representation becomes the Choi matrix of the map and the Fourier inversion formula becomes the Choi inversion formula. We have also derived the Plancherel formula in the context of finite inverse semigroup. Then, by defining the notion of matrix valued positive definite maps, we have proved Bochner’s theorem in the context of finite inverse semigroup. Also, we have proved Schur’s orthogonality relation in the context of finite inverse semigroup which played a crucial role in the proof of Bochner’s theorem. Then we have demonstrated how Bochner’s theorem reduces to Choi theorem on completely positive maps when we consider the inverse semigroup of matrix units. Finally, we have derived the necessary and sufficient condition on a representation $\rho:M_{m}\to M_{d_{\rho}}(\mathbb{C})$ such that the Complete positivity vs. positivity correspondence holds between a linear map $\Phi:M_{m}(\mathbb{C})\to M_{n}(\mathbb{C})$ and its Fourier Transform $\widehat{\Phi}(\rho)$.
@@ Appendix A Proof of Plancherel formula for finite groups
The Plancherel formula can be derived as follows:
$\sum_{g\in G}\Phi(g^{-1})\Psi(g)$ $=\frac{1}{|G|}\sum_{g\in G}\sum_{\rho\in Irr(G)}d_{\rho}\operatorname{tr}_{1}\big{[}(\rho(g)\otimes\mathbb{I})\widehat{\Phi}(\rho)\big{]}\Psi(g)$
$=\frac{1}{|G|}\sum_{g\in G}\sum_{\rho\in Irr(G)}d_{\rho}\operatorname{tr}_{1}\big{[}(\rho(g)\otimes\mathbb{I})\widehat{\Phi}(\rho)(\mathbb{I}\otimes\Psi(g))\big{]}$
$=\frac{1}{|G|}\sum_{g\in G}\sum_{\rho\in Irr(G)}d_{\rho}\operatorname{tr}_{1}\big{[}\widehat{\Phi}(\rho)(\mathbb{I}\otimes\Psi(g))(\rho(g)\otimes\mathbb{I})\big{]}$
$=\frac{1}{|G|}\sum_{g\in G}\sum_{\rho\in Irr(G)}d_{\rho}\operatorname{tr}_{1}\big{[}\widehat{\Phi}(\rho)(\rho(g)\otimes\Psi(g))\big{]}$
$=\frac{1}{|G|}\sum_{g\in G}\sum_{\rho\in Irr(G)}d_{\rho}\operatorname{tr}_{1}\big{[}\widehat{\Phi}(\rho)\widehat{\Psi}(\rho)\big{]}.$
@@
Appendix B Proof of Theorem (A)
@@
Now in the group multiplication table, along a row or column, each element of the group appears once. So it is guaranteed that the terms $g_{i}g_{j}$ for all possible values of $i$ and $j$ produces all the elements in the group. We collect together all the terms in the summation of Eq. (98) for which $g_{i}g_{j}$ is equal to a particular group element say $g_{k}$ and rewrite the summation as follows:
@@
Appendix C Proof of Proposition( 6)
Proposition 6.
For a finite inverse semigroup $S$ and $z\neq s,t\in S$, let $ss^{-1}=tt^{-1}$. Then $s^{-1}t$ is an idempotent if and only if $s=t$.
Proof: When $s=t$, trivially $s^{-1}t$ is an idempotent.
Now, let $s^{-1}t$ is an idempotent. Then
$s^{-1}ts^{-1}t=s^{-1}t$ (102)
$s^{-1}t=t^{-1}s$ (103)
$(s^{-1}t)^{-1}(s^{-1}t)=(s^{-1}t)(s^{-1}t)^{-1}.$ (104)
Using Eq. (104) and the assumption $ss^{-1}=tt^{-1}$, we have
$s^{-1}s=t^{-1}t.$ (105)
Now we use Eq. (103) in Eq. (102) to obtain
$s^{-1}tt^{-1}s=s^{-1}t$
$\iff s^{-1}ss^{-1}s=s^{-1}t$
$\iff s^{-1}s=s^{-1}t$
$\iff ts^{-1}s=ts^{-1}t$
$\iff tt^{-1}t=ts^{-1}t$ (106)
$\iff t=ts^{-1}t,$ (107)
where in Eq. (106) we have used eq. (105). Since in a finite inverse semigroup every element has a unique inverse, we conclude from Eq. (107) that $s=t$. This completes the proof. $\Box$
Appendix D Proof of Bochner’s theorem for matrix valued maps on finite group
We recall that the inversion formula [see Eq. (33)] for maps on finite group is given by
$\Phi(g)=\frac{1}{|G|}\sum_{\rho\in\mathrm{Irr}(G)}d_{\rho}\operatorname{tr}_{d_{\rho}}\big{[}(\rho(g^{-1})\otimes\mathbb{I})\widehat{\Phi}(\rho)\big{]},$
$\Phi(g^{-1}g^{\prime})=\frac{1}{|G|}\sum_{\rho\in\mathrm{Irr}(G)}d_{\rho}\operatorname{tr}_{d_{\rho}}\big{[}(\rho(g^{\prime-1}g)\otimes\mathbb{I})\widehat{\Phi}(\rho)\big{]}.$
Now, using the above equation we have
$\sum_{g,g^{\prime}\in G}\overline{\left<\Phi(g^{-1}g^{\prime})h_{g^{\prime}},h_{g}\right>}$ $=\sum_{g,g^{\prime}\in G}\frac{1}{|G|}\sum_{\rho\in\mathrm{Irr}(G)}d_{\rho}\overline{\left<\operatorname{tr}_{d_{\rho}}\big{[}(\rho(g^{\prime-1}g)\otimes\mathbb{I})\widehat{\Phi}(\rho)\big{]}h_{g^{\prime}},h_{g}\right>}$
@@ Appendix D Proof of Bochner’s theorem for matrix valued maps on finite group
$=\frac{1}{|G|}\sum_{\rho\in\mathrm{Irr}(G)}d_{\rho}\mathrm{tr}\left(\left(\sum_{g^{\prime}\in G}\rho(g^{\prime})^{\dagger}\otimes h_{g^{\prime}}\right)\left(\sum_{g\in G}\rho(g)\otimes h_{g}^{\dagger}\right)\widehat{\Phi}(\rho)\right)$
@@
where $A(\rho,h):=\sum_{g\in G}\rho(g)\otimes h_{g}^{\dagger}$ and $h:=\{h_{g}\}_{g\in G}$ Now, if we assume that $\widehat{\Phi}(\rho)$ is positive semidefinite for all $\rho\in\mathrm{IrrU}(G)$, then we have $\sum_{g,g^{\prime}\in G}\overline{\left<\Phi(g^{-1}g^{\prime})h_{g^{\prime}},h_{g}\right>}\geq 0$ for all subset $h\subset\mathbb{C}^{n}$ implying that $\Phi$ is a positive definite map. This completes the proof in one direction.
@@
where $B\in M_{d_{\rho^{0}}}(\mathbb{C})\otimes\mathbb{C}^{n}$ and $\operatorname{tr}_{d_{\rho^{0}}}$ represents trace over the space $M_{d_{\rho^{0}}}(\mathbb{C})$. Now, using Schur orthogonality relations, we have
@@
Now by substituting the above form of $A(\rho,h^{0})$ in Eq. (109), we get
@@
Now if we assume that $\Phi:G\to M_{n}(\mathbb{C})$ is positive definite then the left hand side of the above equation is non-negative. Since, $B$ and $\rho^{0}$ are chosen arbitrarily, $\widehat{\Phi}(\rho)$ is positive semidefinite for all $\rho\in\mathrm{IrrU}(G)$. This completes the proof.

### v2

@@ 3.5 Representations Characterizing Complete Positivity via Fourier Positivity
Let $\Phi:M_{m}(\mathbb{C})\to M_{n}(\mathbb{C})$ be a linear map and let $\rho:M_{m}\to M_{d_{\rho}}(\mathbb{C})$ be a representation. The correspondence between the complete positivity of $\Phi$ and the positivity of its Fourier transform $\widehat{\Phi}(\rho)$ holds if and only if $\rho$ is of the form $\rho(X)=UXU^{\dagger}$ for all $X\in M_{m}(\mathbb{C})$, where $U:\mathbb{C}^{m}\to\mathbb{C}^{d_{\rho}}$ is unitary with $d_{\rho}=m$.
@@
$=(\mathrm{id}\otimes\Phi)\sum_{i,j}\rho(e_{ij})\otimes e_{ij}.$ (114)
Now it can be shown that
$\sum_{i,j}\rho(e_{ij})\otimes e_{ij}=\tau(C_{\rho^{*}}),$
where $\rho^{*}$ is the adjoint of the representation $\rho:M_{m}(\mathbb{C})\to M_{d_{\rho}}(\mathbb{C})$, $C_{\rho^{*}}$ is the Choi matrix of $\rho^{*}$, and $\tau$ denotes matrix transposition. Consequently,
$\widehat{\Phi}(\rho)$ $=(\mathrm{id}\otimes\Phi)\tau(C_{\rho^{*}})=(\tau\otimes\Phi\circ\tau)C_{\rho^{*}}$ (115)
which is equivalent to
$\tau(\widehat{\Phi}(\rho))=(\mathrm{id}\otimes\tau\circ\Phi\circ\tau)C_{\rho^{*}}.$ (116)
@@
Observe that $\tau(\widehat{\Phi}(\rho))$ is positive semidefinite if and only if $\widehat{\Phi}(\rho)$ is positive semidefinite, and $\Phi$ is CP if and only if $\tau\circ\Phi\circ\tau$ is CP. Now, by Theorem M, the CP-positivity correspondence holds if and only if $\rho^{*}$ is a complete order isomorphism, equivalently if and only if $\rho$ is a complete order isomorphism. Since $\rho$ is a representation, it must be of the form $\rho(X)=UXU^{\dagger}$ for all $X\in M_{m}(\mathbb{C})$, where $U:\mathbb{C}^{m}\to\mathbb{C}^{d_{\rho}}$ is unitary with $d_{\rho}=m$. This completes the proof. $\Box$
4 Conclusion
Bochner’s theorem, one of the most celebrated theorems in harmonic analysis, provides a precise correspondence between positive definite maps and the positivity of their Fourier transforms. The central contribution of this work is the extension of this correspondence to the setting of finite inverse semigroups.
We establish a Bochner-type theorem for matrix-valued maps on the contracted algebra of a finite inverse semigroup. Exploiting the intrinsic partial order of inverse semigroups, positivity naturally arises at the level of Möbius transformed map. Our main result characterizes the positive definiteness of this transformed map in terms of the positivity of the Fourier transform of the original map with respect to a complete family of inequivalent irreducible representations of the contracted algebra induced by a complete family of inequivalent irreducible unitary representations of the maximal subgroups of the inverse semigroup.
The proof of this Bochner-type theorem is supported by several structural results developed here for finite inverse semigroups. These include a convolution algebra structure on the space of linear maps, a Fourier transform and inversion formula, a Plancherel formula, alternative characterizations of positive definite maps, and a Schur orthogonality relation adapted to the inverse semigroup setting.
Within this framework, Choi’s theorem on completely positive maps arises naturally as a special case when the inverse semigroup of matrix units are considered. Moreover, the Fourier transform with respect to identity representation coincides with the Choi matrix of the linear map, and the Fourier inversion formula reduces to the Choi inversion formula. This shows that Choi’s characterization of complete positivity is a manifestation of a more general Bochner-type characterization of positive definite maps.
@@ Appendix A Proof of Plancherel Formula for Finite Groups
Let $G$ be a finite group, and let $\Phi,\Psi:G\to M_{n}(\mathbb{C})$ be linear maps. Using the Fourier inversion formula for maps on $G$, we have
$\Phi(g^{-1})$ $=\frac{1}{|G|}\sum_{\rho\in\mathrm{Irr}(G)}d_{\rho}\operatorname{tr}_{d_{\rho}}\big[(\rho(g)\otimes\mathbb{I})\widehat{\Phi}(\rho)\big].$ (117)
Substituting this expression into $\sum_{g\in G}\Phi(g^{-1})\Psi(g)$ and rearranging the terms yields
$\sum_{g\in G}\Phi(g^{-1})\Psi(g)$ $=\frac{1}{|G|}\sum_{g\in G}\sum_{\rho\in\mathrm{Irr}(G)}d_{\rho}\operatorname{tr}_{d_{\rho}}\big[(\rho(g)\otimes\mathbb{I})\widehat{\Phi}(\rho)(\mathbb{I}\otimes\Psi(g))\big]$
$=\frac{1}{|G|}\sum_{g\in G}\sum_{\rho\in\mathrm{Irr}(G)}d_{\rho}\operatorname{tr}_{d_{\rho}}\big[\widehat{\Phi}(\rho)(\mathbb{I}\otimes\Psi(g))(\rho(g)\otimes\mathbb{I})\big]$
$=\frac{1}{|G|}\sum_{g\in G}\sum_{\rho\in\mathrm{Irr}(G)}d_{\rho}\operatorname{tr}_{d_{\rho}}\big[\widehat{\Phi}(\rho)(\rho(g)\otimes\Psi(g))\big]$
$=\frac{1}{|G|}\sum_{\rho\in\mathrm{Irr}(G)}d_{\rho}\operatorname{tr}_{d_{\rho}}\big[\widehat{\Phi}(\rho)\widehat{\Psi}(\rho)\big].$
@@
Appendix B Proof of Theorem A
@@
Now, in the group multiplication table, along a row or column, each element of the group appears once. So, it is guaranteed that the terms $g_{i}g_{j}$ for all possible values of $i$ and $j$ produce all elements of the group. We collect together all the terms in the summation of Eq. (118) for which $g_{i}g_{j}$ is equal to a particular group element, say $g_{k}$ and rewrite the summation as follows:
@@
Appendix C Proof of Bochner’s theorem for matrix valued maps on finite group
Let $G$ be a finite group, and let $\Phi:G\to M_{n}(\mathbb{C})$ be a linear map. Recall the Fourier inversion formula [see Eq. (40)] for maps on finite group:
$\Phi(g)=\frac{1}{|G|}\sum_{\rho\in\mathrm{Irr}(G)}d_{\rho}\operatorname{tr}_{d_{\rho}}\big[(\rho(g^{-1})\otimes\mathbb{I})\widehat{\Phi}(\rho)\big].$ (122)
Consequently,
$\Phi(g^{-1}g^{\prime})=\frac{1}{|G|}\sum_{\rho\in\mathrm{Irr}(G)}d_{\rho}\operatorname{tr}_{d_{\rho}}\big[(\rho(g^{\prime-1}g)\otimes\mathbb{I})\widehat{\Phi}(\rho)\big].$ (123)
Now, using the above equation, we compute
$\sum_{g,g^{\prime}\in G}\overline{\left<\Phi(g^{-1}g^{\prime})h_{g^{\prime}},h_{g}\right>}$
$=\sum_{g,g^{\prime}\in G}\frac{1}{|G|}\sum_{\rho\in\mathrm{Irr}(G)}d_{\rho}\overline{\left<\operatorname{tr}_{d_{\rho}}\big[(\rho(g^{\prime-1}g)\otimes\mathbb{I})\widehat{\Phi}(\rho)\big]h_{g^{\prime}},h_{g}\right>}$
@@ Appendix C Proof of Bochner’s theorem for matrix valued maps on finite group
$=\frac{1}{|G|}\sum_{\rho\in\mathrm{Irr}(G)}d_{\rho}\mathrm{tr}\left(\Bigg(\sum_{g^{\prime}\in G}\rho(g^{\prime})^{\dagger}\otimes h_{g^{\prime}}\Bigg)\Bigg(\sum_{g\in G}\rho(g)\otimes h_{g}^{\dagger}\Bigg)\widehat{\Phi}(\rho)\right)$
@@
where
$A(\rho,h):=\sum_{g\in G}\rho(g)\otimes h_{g}^{\dagger}$
and $h:=\{h_{g}\}_{g\in G}\subset\mathbb{C}^{n}$.
Now, if we assume that $\widehat{\Phi}(\rho)$ is positive semidefinite for all $\rho\in\mathrm{IrrU}(G)$, then from Eq. (124) we have $\sum_{g,g^{\prime}\in G}\overline{\left<\Phi(g^{-1}g^{\prime})h_{g^{\prime}},h_{g}\right>}\geq 0$ for all subsets $h\subset\mathbb{C}^{n}$, which implies that $\Phi$ is a positive definite map. This completes the proof in one direction.
@@
where $B\in M_{d_{\rho^{0}}}(\mathbb{C})\otimes\mathbb{C}^{n*}$, and $\operatorname{tr}_{d_{\rho^{0}}}$ denote the trace over the space $M_{d_{\rho^{0}}}(\mathbb{C})$. Now, using Schur orthogonality relations, we have
@@
Consequently, substituting the above form of $A(\rho,h^{0})$ in Eq. (124), we obtain
@@
Now if we assume that $\Phi:G\to M_{n}(\mathbb{C})$ is positive definite, then the left hand side of the above equation is non-negative. Since $B$ and $\rho^{0}$ are chosen arbitrarily, $\widehat{\Phi}(\rho)$ is positive semidefinite for all $\rho\in\mathrm{IrrU}(G)$. This completes the proof.